A simple pendulum has a time period T1 when on the earths surface and T2 when taken to height R above the earths surface where R is the radius of the earth the value of T1 /T2 is

Grade:11

2 Answers

Arun
6 years ago
Dear Student
 
 
T=2pi*sqrt(l/g)
g is inversly proportional to (R+h)2 …..........h is height above the earth’s  surface
therefore T is proportional to R+h
T2/T1= (R+h)/R =2............ putting h=R
Hence 
T1 / T2  = 1/2 RegardsArun (askIITians forum expert)
bharat
6 years ago
time period of a pendulum on earth is 
T=2\prod*sqrt(r/g)
g is the acceleration due to gravity
as now the distance is r+h 
since g is inversely proportional to (r+h)^2  
where h is the height above the earth’s surface
hence t is proportinal to r+h
now since h=r
t2/t1=(r+h)/r=2 
hence 
t1/t2=1/2